# Towards Scalable Exact Machine Unlearning Using Parameter-Efficient Fine-Tuning

### v1

@@ 3.3 Training under Budget Constraints
The training strategy discussed in the previous section requires training the same model on $L!$ slice sequences, which is expensive in practice. In this section, we discuss strategies to select sequences under a budget, $B$ (maximum number of sequences that can trained). First, we show that there exists an optimal subset of $B$ slice sequences and randomly selecting $B$ sequences may not be effective. To illustrate this idea, we use a permutation tree [4], where all possible permutation sequences are embedded into a tree structure. In Figure 3 (left), we show an example of a permutation tree with $L=3$ slices, paths from the root to the leaves correspond to unique slice sequences $(S_{1},S_{2},S_{3})$, $(S_{1},S_{3},S_{2})$, and so on. From the previous section, we know that the topmost slice is most sensitive because if a deletion request affects the topmost slice the entire model needs to be retrained (shown in Figure 2 (right)). To address this and reduce the retraining cost, we should ensure we train models on sequences with different top slices. Building on this intuition, in the general setting we should train the model on diverse permutation sequences. An example is illustrated in Figure 3 (left), where a diverse set of 3 sequences is marked in green (where no identical slices occupy the same position). Selecting diverse permutations is challenging for two reasons: (a) relying solely on random sampling may not always yield optimal results, and (b) performing naive random sampling may not be feasible with large slice counts, e.g., $L=32$, where even enumerating all 32! permutations becomes prohibitively expensive. Next, we provide two strategies to select permutation sequences for a budget $B$.
@@
Uniform Deletion Prior, Budget $B\leq L$. The first strategy is called Bipartite-matching based Selection (BMS), which is designed for the setting where budget $B\leq L$ and each slice has a uniform probability of being impacted by deletion. This scenario is quite common when the budget is limited and deletion probabilities for slices are unknown. In this setting, finding the most diverse permutation set can be decomposed into solving the perfect matching in bipartite graphs [21]. We describe the intuition behind BMS using a permutation tree in Figure 3 (center), wherein we sequentially select elements of the permutations. The exact functioning of BMS is shown in Figure 4. BMS iteratively selects elements of sequences by defining a bipartite graph between one level to the next one using the edges of the permutation tree (where edges are incident on feasible elements that haven’t been encountered in the sequence). Selecting the next element then is equivalent to finding a perfect matching [19] between one level and the next one as shown by the bold lines in Figure 4. This continues till all sequences have $L$ elements. Overall, BMS has a time complexity of $O(L^{2.5})$ (details in Appendix B.1). The following result shows that BMS is optimal (proof in Appendix A.1).
Lemma 1.
Known Deletion Prior. The next strategy addresses the general case where $B$ can be anything and assume access to deletion probabilities for each slice. For simplicity, we consider the probability of a deletion request affecting a slice as $\{p_{1},\ldots,p_{L}\}$. The intuition behind this approach is to ensure that slices with a lower probability of deletion are placed at the start of the sequence. Therefore, we use a conditional sampling strategy to construct individual sequences by sampling one slice at a time according to the probability of not being deleted ($1-p_{i}$). This process is shown in Figure 3 (right), where we select the next slice in a sequence. We repeat this sampling process till we have $B$ unique sequences and use them for training. We can use the same algorithm when deletion priors are unknown by considering a uniform distribution. Note that conditional sampling may not return the optimal permutation subset and has similar guarantees as random sampling but without its computational bottleneck. Selecting the $B$ most diverse permutations is a challenging combinatorial problem, which to the best of our knowledge, has not been extensively studied in the algorithms literature. We defer the task of developing a general algorithm for this problem to future works.
@@ 3.4 Theoretical Analysis
In this section, we theoretically analyze the performance of exact unlearning system. For this, we draw inspiration from approximate unlearning literature [51] and introduce the definition of deletion rate for exact unlearning systems.
@@
The proof is presented in Appendix A.2. The above result is based on the simple observation that when the budget $B\geq L$, deletion requests should impact all $mL$ slices ($m$: number of shards, $L$: slices per shard) across shards for S3T to encounter service failure. However, in the case of SISA, if deletion requests affect the first slice of every shard (total of $m$ slices), then the system will experience service failure and need to be retrained from scratch.
Next, we analyze the impact of deletion requests on the system’s performance. We perform a fine-grained analysis focusing on an individual shard performance. In this setting, we consider the real-world scenario where we do not retrain every time a slice is impacted instead work with the best available model. For S3T, this means switching off the necessary PEFT layers, while for SISA, it means reverting to the best available checkpoint. The unlearning system experiences performance degradation with an increasing number of deletion requests, as we are compelled to utilize a model trained on fewer data slices (we show this empirically in Section 4.2). To quantify the performance retention we use a monotonically increasing function $F(k)$, which indicates a model’s performance when trained on $k$ slices. The exact formulation of $F(\cdot)$ depends on several factors like the dataset, model size, etc. We analyze the performance retention while processing deletion requests.
@@
Given a budget $B$, the probability that a shard retains a performance, $F_{r}(\cdot)$, of at least $F(k)$ after $r$ deletion requests is shown below
@@
For SISA, the same performance guarantee is $\mathbb{P}\left[F_{r}(\mathrm{SISA})\geq F(k)\right]=(1-k/L)^{r}$.
The proof is presented in Appendix A.3. The above result demonstrates that compared to SISA S3T significantly enhances performance by increasing the probability that the system maintains at least $F(k)$ performance by a factor of $B^{\prime}$. This shows that increasing the training budget $B$ helps improve the performance of the unlearning system. Another takeaway from the above lemma is that to ensure the system performance of $F(k)$, there is no point in increasing the budget $B$ beyond $\binom{L}{k}$.
@@ 4.1 Sequential Slice-wise Training Performance
In Figure 5, we report the performance of fine-tuning on vision, GLUE, and SuperGLUE benchmarks. We use ViTBASE [15] (for CIFAR10 & CIFAR100 [34]), ViTLARGE (for Tiny ImageNet [38]), and RoBERTaLARGE [42] (for GLUE [62] & SuperGLUE [61]). We observe that S3T achieves comparable performance to full training (FT) across all settings. In some of the settings, we also observe that S3T is able to outperform FT (e.g., S3T obtains 2.5% accuracy gain on TinyImagenet using ViT-large). Next, we conduct experiments to evaluate the effectiveness of S3T while using large language models (LLMs). Specifically, we perform instruction tuning of Llama2-7B [60], Llama2-13B, and Llama3-8B using Alpaca dataset [55]. Then, we evaluate each instruction-tuned model on a range of tasks to evaluate the model’s general/world knowledge (MMLU [29], OpenBookQA [44]), truthfulness in QA (TruthfulQA [39]), and commonsense reasoning (PIQA [6], HellaSwag [67], Winogrande [50], ARC [12]). We use LLM-evaluation suite [20] to report the performance and report the zero-shot performance for all datasets. Similar to the previous setup, in Table 1, we observe that S3T achieves comparable performance to full finetuning across all datasets and even outperforms FT on many datasets across different model sizes. These experiments provide the answer to (RQ1) demonstrating that S3T is an effective way to fine-tune Transformer models, enabling us to achieve parameter isolation for data slices without impacting the model’s performance.
@@ 4.2 Deletion Performance
In this section, we evaluate the performance of S3T as deletion requests are received. In Figure 6 (left), we report the performance of S3T and baseline SISA (with $m=5$ shards, $L=4$ slices) on CIFAR-10 and CIFAR-100 datasets. We report S3T’s performance under various budgets. We observe that S3T can achieve very close performance to the full budget ($B=24$) with a significantly smaller budget, $B=4$. We also observe that S3T can handle more deletion requests and consistently outperforms the performance of the baseline method, SISA.
Next, we extensively evaluate the impact of increasing the training budget $B$ on the deletion rate. In Figure 6 (right), we observe that there is a steady growth in the deletion rate with an increasing budget. The growth is slightly higher in the initial stages when the budget is low and slows down gradually. This experiment provides empirical evidence to our theoretical result in Lemma 3, which claims that the performance improves with an increased budget.
@@ 4.3 Analysis
In this section, we conduct analysis experiments to evaluate different components within S3T. [^2] Apart from these experiments, we also report several other analysis experiments in Appendix C.2.
Figure 6: (Left) We report the impact on performance of S3T and baselines with an increasing number of deletion requests. We observe that S3T can handle a higher number of deletion requests while maintaining high performance with a relatively low budget. (Right) We report the deletion rate of S3T with an increasing budget and observe steady growth.
Figure 7: We report the deletion rate of S3T and compare it with the baselines. (Left) We compare the deletion rates of S3T under varying budgets with SISA and observe significant gains. (Center) We report the deletion rates with increasing number of shards and (Right) an increasing number of slices. In both scenarios, we observe that the deletion rate of S3T grows significantly faster than the baseline.
Deletion Ablations. We evaluate the deletion capabilities of S3T and compare with the theoretical bounds. In Figure 7 (left), we report the deletion rate of S3T and SISA for the setup (with $m=5$ shards, $L=32$ slices). We observe that with a small budget $B=8$, S3T can achieve 1.6x gains in the number of deletion requests handled. We also plot the theoretical bounds derived in Section 3.4 and show that they hold in practice. Note that increasing the budget $B>L$ does not help improve the deletion rate but increases the probability of a better-performing model (as observed in Lemma 3).
Figure 8: We evaluate the performance of bipartite matching-based selection (BMS) and observe that BMS consistently outperforms random sampling.
@@
This shows that by using additional offline training (increased budget) it is possible to improve the deletion rate thereby reducing retraining costs.
Sequence Selection. We evaluate the quality of the sequences generated by the BMS algorithm (Section 3.3) and compare it with random sampling. Ideally, we want the selected sequences to be diverse and thereby compute the edit distance between sequences. Specifically, we report the following score for a selected subset, $O$: $Q(O)=\mathbb{E}_{o\in O}[\min_{o^{\prime}\in O\setminus o}d_{\mathrm{edit}}(o,o^{\prime})]$, which measures the average minimum edit distance within the subset. In Figure 8 (left), we report the $Q(O)$ for a varying number of slices, $L$. We observe that BMS consistently outperforms random sampling, which indicates that the generated sequences are more diverse.
@@ 5 Conclusion
In this paper, we introduced S3T, an effective framework for performing exact unlearning. S3T uses a modular machine learning model that is trained using disjoint shards of the data. Each shard is used to train a different model using a lightweight fine-tuning approach that enables parameter isolation, which allows us to execute unlearning requests efficiently. The key idea behind S3T is to train multiple models using different sequences of slices before deploying the system. This helps reduce retraining costs and improve the model’s performance while the model is in production. Both theoretically and empirically, we show that S3T has significantly improved deletion capabilities compared to existing approaches. Future work can focus on developing techniques for finer-grained parameter isolation to further improve the overall performance of S3T.
6 Acknowledgments
The authors thank Vinayshekhar Bannihatti Kumar, Nicholas Monath, and Shounak Chattopadhyay for helpful feedback and discussions.
7 Author Contributions
SBRC proposed the unlearning idea, conducted all the experiments and led the project. KC proposed the diverse permutation selection algorithm using bipartite graph matching. AS contributed to deriving the theoretical results and designing the empirical studies. AD and SC provided ideas and served as project advisors, offering extensive assistance in organizing the manuscript. All authors contributed to the writing of the manuscript.

### v2

@@ 3.3 Training under Budget Constraints
The training strategy discussed in the previous section requires training the same model on $L!$ slice sequences, which is expensive in practice. In this section, we discuss strategies to select sequences under a budget, $B$ (maximum number of sequences that can trained). First, we show that there exists an optimal subset of $B$ slice sequences and randomly selecting $B$ sequences may not be effective. To illustrate this idea, we use a permutation tree (Bhattacharya, 1994), where all possible permutation sequences are embedded into a tree structure.
In Figure 3 (left), we show an example of a permutation tree with $L=3$ slices, paths from the root to the leaves correspond to unique permutation sequences $(S_{1},S_{2},S_{3})$, $(S_{1},S_{3},S_{2})$, and so on. From the previous section, we know that the topmost slice is the most sensitive because if a deletion request affects the topmost slice the entire model needs to be retrained (shown in Figure 2 (right)). To address this and reduce the retraining cost, we should ensure we train models on sequences with different top slices. Building on this intuition, in the general setting we should train the model on diverse permutation sequences. Two sequences are considered diverse if no element appears at the same position, e.g., $(S_{1},S_{2},S_{3})$ and $(S_{2},S_{3},S_{1})$. An example is illustrated in Figure 3 (left), where a diverse set of 3 sequences is marked in green (where no identical slices occupy the same position). Selecting diverse permutations is challenging as relying solely on random sampling may not always yield the best results. Moreover, in certain scenarios, it is possible to have prior knowledge about the deletion probabilities of data slices; for example, younger users might be more likely to request data deletion than older users. Therefore, we present two strategies for selecting diverse permutation sequences for a budget $B$, depending on whether or not prior deletion probabilities are available.
@@
Uniform Deletion Prior. In the setting, where each slice has a uniform (or unknown) deletion prior, we can generate diverse sequences by using cyclic permutations of the original sequence. Given a sequence $(S_{1},S_{2},S_{3})$, the cyclic permutations are (shown in Figure 3 (middle)):
$(S_{1},S_{2},S_{3})\rightarrow(S_{3},S_{1},S_{2})\rightarrow(S_{2},S_{3},S_{1}).$ (2)
The above approach guarantees that no element is repeated in the same position. However, it can only generate up to $L$ different sequences. For budget $B>L$, we extend the cyclic rotation algorithm to generate more sequences. In Figure 3 (right), we generate new sequences by iterating over the existing sequences and performing cyclic rotations of the subsequences. For example, for $(S_{1},S_{2},S_{3})$ we perform cyclic rotation of the 2nd and 3rd element to obtain the sequence: $(S_{1},S_{3},S_{2})$ (more examples in Figure 10). We provide the general iterative cyclic rotation algorithm in Appendix B.1.
Non-uniform Deletion Prior. In scenarios, where we have prior knowledge of the deletion probabilities, sequences generated by cyclic rotation may not be ideal. For example, consider the deletion probabilities are: ($S_{1}$: 0.5, $S_{2}$: 0.4, $S_{3}$: 0.1). Then, the first sequence in Eq. 2 is a bad choice because two of the slices most likely to be deleted are placed at the top. It is possible to select better sequences while satisfying the diversity criteria (no repeated slices at the same position). Before describing the selection algorithm, we show how to evaluate the quality of a sequence. We evaluate a sequence with deletion probabilities: $\mathcal{S}=(p_{1},\ldots,p_{L})$ by computing the expected number of functioning slices after $t$ deletions: $\mathrm{score}[\mathcal{S},t]=\sum_{i=1}^{L}i.\left(1-\sum_{j=1}^{i}p_{j}\right)^{t}$, where $t$ is a hyperparameter.
We present a bipartite-matching based sequence (BMS) selection algorithm. We will provide an intuitive explanation of the algorithm here and refer to Appendix B.2 for complete details. An illustration of BMS is shown in Figure 4. BMS iteratively selects elements of sequences by constructing a bipartite graph between one level to the next one (where edges are incident on feasible elements for the current sequence). The edges are weighted by the score of the current sequence, $\mathrm{score}[\mathcal{S},t]$. Selecting the next element then is equivalent to finding a maximum weight perfect matching (Galil, 1986) using the Hungarian algorithm (Kuhn, 1955) shown by the bold lines in Figure 4. This continues till all sequences have $L$ elements. The BMS algorithm can generate up to $L$ sequences. For a budget $B>L$, we use conditional sampling to randomly generate sequences according to their deletion probabilities (see details in Appendix B.2).
Overall Workflow. So far, we have discussed the sequential training procedure and the sequence selection approach within a budget. Here, we will bring all of the components together and illustrate the functioning of S3T using a simple example. We consider a setting with $m=3$ shards, $L=4$ slices, and budget $B=4$. Initially, for every shard, the available models are trained on sequences:
$(1,2,3,4),(4,1,2,3),(3,4,1,2),(2,3,4,1)$
Next, if a deletion request affects slice 1 for the 3rd shard. Then, the available models for the 3rd shard are: $(4),(3,4),(2,3,4)$. Notice how all the PEFT layers at or below slice 1 have been switched off. In this scenario, S3T doesn’t perform any retraining but continues to function with the best available model: $(2,3,4)$ (as it is trained on maximum slices). If the following deletion requests affect slice 2, then the available models are: $(4),(3,4)$. S3T continues to function with the best model $(3,4)$. This continues until no models are available in any shard, at which point S3T is retrained from scratch.
@@ 3.4 Theoretical Analysis
In this section, we theoretically analyze the performance of exact unlearning system. For this, we draw inspiration from approximate unlearning literature (Sekhari et al., 2021) and introduce the definition of deletion rate for exact unlearning systems.
@@
The above result states that the deletion rate doesn’t improve by increasing the budget $B$ beyond $L$. This shows that the optimal deletion rate can be achieved using only $L$ sequences (instead of $L!$). The complete proof is presented in Appendix A.1.
Next, we analyze the impact of deletion requests on the system’s performance. We perform a fine-grained analysis focusing on the performance of an individual shard. In this setting, we consider the real-world scenario where we do not retrain every time a slice is impacted instead work with the best available model. For S3T, this means switching off the necessary PEFT layers, while for SISA, it means reverting to the best available checkpoint. The unlearning system experiences performance degradation with an increasing number of deletion requests, as we are compelled to utilize a model trained on fewer data slices (we show this empirically in Section 4.2). To quantify the performance retention we use a monotonically increasing function $F(k)$, which indicates a model’s performance when trained on $k$ slices. The exact formulation of $F(\cdot)$ depends on several factors like the dataset, model size, etc. We analyze the performance retention while processing deletion requests.
@@
Given a set of randomly selected $B>1$ sequences and uniform deletion prior of slices, the difference in the probability that a shard retains a performance, $F_{r}(\cdot)$ of at least $F(k)$ after $r$ deletion requests, between S3T and SISA, is shown below
@@
where $\zeta=1-(1-k/L)^{r}$ is a positive fraction and $B^{\prime}=\min\left\{B,\frac{L!}{(L-k)!}\right\}$.
The above result demonstrates that compared to SISA, S3T significantly enhances performance by increasing the probability that the system maintains at least $F(k)$ performance. Eq. 3 shows that increasing the training budget $B$ helps improve the performance of the unlearning system. The proof is presented in Appendix A.2. Another takeaway from the above lemma is that to ensure the system performance of $F(k)$, there is no point in increasing the budget $B$ beyond $\frac{L!}{(L-k)!}$.
@@ 4.1 Sequential Slice-wise Training Performance
In Figure 5, we report the performance of fine-tuning on vision, GLUE, and SuperGLUE benchmarks. We use ViTBASE (Dosovitskiy et al., 2020) (for CIFAR10 & CIFAR100 (Krizhevsky et al., 2009)), ViTLARGE (for Tiny ImageNet (Le & Yang, 2015)), and RoBERTaLARGE (Liu et al., 2019) (for GLUE (Wang et al., 2018) & SuperGLUE (Wang et al., 2019)). We observe that S3T achieves comparable performance to full training (FT) across all settings. In some of the settings, we also observe that S3T is able to outperform FT (e.g., S3T obtains 2.5% accuracy gain on TinyImagenet using ViT-large). Next, we conduct experiments to evaluate the effectiveness of S3T while using large language models (LLMs). Specifically, we perform instruction tuning of Llama2-7B (Touvron et al., 2023), Llama2-13B, and Llama3-8B using Alpaca dataset (Taori et al., 2023). Then, we evaluate each instruction-tuned model on a range of tasks to evaluate the model’s general/world knowledge (MMLU (Hendrycks et al., 2020), OpenBookQA (Mihaylov et al., 2018)), truthfulness in QA (TruthfulQA (Lin et al., 2022)), and commonsense reasoning (PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), ARC (Clark et al., 2018)). We use LLM-evaluation suite (Gao et al., 2023) report the zero-shot performance for all datasets. Similar to the previous setup, in Table 1, we observe that S3T achieves comparable performance to full finetuning across all datasets and even outperforms FT on many datasets across different model sizes. These experiments provide the answer to (RQ1) demonstrating that S3T is an effective way to fine-tune Transformer models, enabling us to achieve parameter isolation for data slices without impacting the model’s performance.
@@ 4.2 Deletion Performance
In this section, we evaluate the performance of S3T as deletion requests are received. In Figure 6 (left), we report the performance of S3T, SISA (Bourtoule et al., 2021), and ProtoSISA (Yan et al., 2022) on CIFAR-10 and CIFAR-100 datasets (all systems use $m=5$ shards, $L=4$ slices). In this experiment, we consider a uniform deletion prior over all slices. We also report the performance of full re-training, which retrains the model after each deletion request and serves as an upper performance bound. We report S3T’s performance under various budgets. We observe that S3T can achieve very close performance to the full budget ($B=24$) with a significantly smaller budget, $B=4$. For SISA and ProtoSISA, we observe that the plot ends after approximately 40 deletion requests as these systems do not have functioning models beyond this point. We observe that S3T can handle more deletion requests while consistently outperforming the baseline approaches. Note that increasing the budget $B>L$ does not help improve the deletion rate but increases the probability of a better-performing model (as observed in Lemma 1).
Next, we extensively evaluate the impact of increasing the training budget $B$ on the deletion rate (with $m=5$ shards & $L=64$ slices). In Figure 6 (right), we observe that there is a steady growth in the deletion rate with an increasing budget. The growth is slightly higher in the initial stages when the budget is low and slows down gradually. This experiment provides empirical evidence to our theoretical result in Lemma 2, which claims that the performance improves with an increased budget.
@@ 4.3 Analysis
In this section, we conduct analysis experiments to evaluate different components within S3T. We also report several other analysis experiments in Appendix C.3.
[…]
Figure 7: We report the deletion rate of S3T and compare it with baselines. (Left) We compare the deletion rates of S3T under varying budgets with SISA and observe significant gains. (Center) We report the deletion rates with an increasing number of shards and (Right) an increasing number of slices. In both scenarios, we observe that S3T’s deletion rate grows significantly faster than SISA.
Deletion Ablations. We evaluate the deletion capabilities of S3T and compare with the theoretical bounds. In Figure 7 (left), we report the deletion rate of S3T and SISA for the setup (with $m=5$ shards, $L=32$ slices). We observe that with a small budget $B=8$, S3T can achieve 1.6x gains in the number of deletion requests handled. We also plot the theoretical bounds derived in Section 3.4 and show that they hold in practice.
@@
Figure 8: We evaluate the performance of iterative cyclic rotation and bipartite matching-based selection (BMS). (Left) We observe that cyclic rotation selection consistently outperforms random sampling for all budgets, $1\leq B\leq 120$ (with fixed $L=5$). (Center) We evaluate the average edit distance of the sequences generated by BMS and observe that it achieves the optimal edit distance ($L$). (Right) We also observe that sequences from BMS achieves higher scores than random sampling.
Figure 9: We compare the deletion time as the number of deletion requests increases. S3T requires the lowest deletion time compared to SISA and full re-training.
Sequence Selection. We evaluate the quality of the sequences generated by the iterative cyclic rotation and BMS algorithm (Section 3.3). Ideally, we want the selected sequences to be diverse and should have a high edit distance between sequences. Therefore, we report the average edit distance within a selected subset, $O$: $\mathbb{E}_{o\in O}[d_{\mathrm{edit}}(o,o^{\prime})]$. First, we consider the setting where the deletion prior is uniform and compare cyclic rotation with random sampling. In Figure 8 (left), we report the average edit distance with varying budget ($B$) while the slice count ($L$) is fixed. We observe that cyclic rotation produces significantly better sequences than random sampling. Second, we consider slices associated with a deletion prior. We present the results by averaging over 10 deletion priors sampled from a Dirichlet distribution. In Figure 8 (center & right), we observe that BMS consistently outperforms random sampling both in terms of diversity (avg. edit distance) and chosen sequence scores ($\mathrm{score}[\mathcal{S},t]$).
Deletion Time. Unlearning in S3T is cost-effective; it primarily involves selecting the best checkpoint and swapping with the current one. The main cost is incurred when a large number of deletion requests necessitate re-training the system from scratch. In Figure 9, we report the total deletion time as the number of deletion requests increases, which includes the checkpoint swapping and re-training time. In this experiment, we compare the deletion time of S3T with SISA and full re-training on CIFAR10 dataset. Full re-training retrains the model after each deletion request. For SISA and S3T, the step jumps indicate that the system requires retraining. Overall, S3T achieves the lowest deletion time. S3T reduces the total deletion time (after 1000 requests) by 2.8x compared to SISA and 25x compared to full retraining. This experiment shows the efficacy of S3T in reducing the overall deletion time.
@@ 5 Conclusion
In this paper, we introduced S3T, an effective framework for performing exact unlearning. S3T uses a modular machine learning model that is trained using disjoint shards of the data. Each shard is used to train a different model using a lightweight fine-tuning approach that enables parameter isolation, which allows us to execute unlearning requests efficiently. The key idea behind S3T is to train multiple models using different sequences of slices before deploying the system. This helps reduce retraining costs and improve the model’s performance while the model is in production. Both theoretically and empirically, we show that S3T has significantly improved deletion capabilities compared to existing approaches. Future work can focus on developing techniques for finer-grained parameter isolation to further improve S3T’s performance.
Acknowledgments
The authors thank Anneliese Brei, Anvesh Rao, Archiki Prasad, Haoyuan Li, Nicholas Monath, Rajeev Ambati, Rakesh Menon, Shounak Chattopadhyay, and Vinayshekhar Bannihatti Kumar for helpful discussions and feedback on the draft.
Author Contributions
SBRC proposed the unlearning idea, conducted all the experiments, and led the project. KC proposed the diverse permutation selection algorithm using bipartite graph matching. AS contributed to deriving the theoretical results and designing the empirical studies. AD and SC provided ideas and served as project advisors, offering extensive assistance in organizing the manuscript. All authors contributed to the writing of the manuscript.